### v1

# Carbon-based light emitting defects in different layers of hexagonal boron nitride

Abstract
Substitutional carbon defects in hexagonal boron nitride (hBN) have garnered significant interest as single photon emitters (SPEs) due to their remarkable optical and quantum properties. An intriguing property of these defects is that they can be spin-active (), even if weakly interacting. Employing density functional theory (DFT) calculations, we demonstrate that two monomers of C-based defects of the same species can exhibit a stable triplet spin state at room temperature, even when they are separated nm, if they reside in different layers. The zero-phonon line (ZPL) energy of C defects in different layers lies within eV range. Also, we found defects that deviate from the typical phonon replica patterns, potentially explaining the observed phonon replicas in yellow emitters in hBN.
I Introduction
Hexagonal boron nitride (hBN) is a layered material that can be exfoliated down to the monolayer. It has garnered attention for hosting single photon emitters [1, 2, 3]. Many of these defects have been associated with substitutional carbon defects [1, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], even though other options have been proposed [14, 15, 16]. Due to its large bandgap, close to 6 eV, defects incorporated within the crystal can introduce new energy levels within the gap [17, 18]. Consequently, electronic transitions give rise to SPEs exhibiting characteristics such as bright photons, a commendable coherence time [19], even at room temperature, and in some cases, they are spin-active (triplet or higher spin state) defects [20, 21, 22, 23]. These properties are crucial for quantum technologies in communication, computation, metrology, sensing, imaging, etc. [24, 25, 26, 27, 28, 29, 30]. Therefore, the search for these types of defects is relevant, considering the scarcity of substitutional spin-active carbon defects, such as CNCN [31], (CN)3CB and (CB)3CN [32, 33]. It is worth mentioning that vacancy-based SPEs in hBN are an active research field, as they often are spin-active [34, 35, 36, 37].
Substitutional C defects behave as donors (CB), acceptors (CN), donor-acceptor pairs, or more complex arrangements [38, 39]. Each C defect contributes a single electron and one defect state (spin up and down) within the hBN band gap. The 4.1 eV emission line has been identified as two adjacent CNCB defects [7], often denoted as C2. If the same CN–CB pair is in the same layer but separated by some distance, the emission energy decreases to the visible range [40, 25]. The ground state is likely spinless if the number of CN and CB is the same. Otherwise, the system can have a non-zero net spin. There are experimental findings of emissions attributed to C defects, even though the actual C defect has not been identified [20, 41]. One possible explanation, fitting with the available results, is weakly coupled defects [42].
This article aims to investigate SPEs when forming a C dimer but residing in different hBN layers. Such a possibility has not been addressed before, even though defects comprising an interstitial C atom have been considered [43]. The paper is structured as follows: Section II details the methods employed for the calculations, Section III presents our main results, including the SPE’s spin state - likely to be a triplet - and their photoluminescence spectra. Finally, Section IV explains the large energy difference between triplet and singlet states for some dimers, the phonon sidebands shape, and we relate our findings to experimental results on spin-active SPEs.
II Methodology
II.1 Computational Methods
We used density functional theory (DFT) as implemented in the Vienna ab initio simulation package (VASP) [44, 45, 46, 47]. A hexagonal supercell was employed together with a single k-point (). We used an hBN bilayer to simulate defects in adjacent layers. If the defects are in non-adjacent layers, we used a hBN trilayer. The connection with the bulk is almost immediate; the additional layers should have (i) a lesser effect on the ZPL energy due to enhanced screening [48, 49, 50] and a shift in the ZPL due to electrostatic potential [51]. We used the AA’ stacking (B, N atoms on top of N, B), as found in most hBN samples [52].
The kinetic energy cutoff was set to 500 eV. PyProcar [53, 54] was used for the analysis of the results, and VESTA for visualization [55].
We used the SCAN exchange-correlation functional [56] to relax the hBN structure and to calculate phonons. Although this functional can provide an accurate description of the geometry, it could fail to accurately describe the defect levels [57]. Once the geometry was obtained with SCAN, we used the Heyd-Scuseria-Ernzerhof (HSE06) functional to obtain the energies and the optical transitions at that geometry [58, 59]. This procedure was used for the ground and excited states. To study the excited state of each configuration, the SCF method was used [60, 61]. This method has a problem of correctly describing the excited state, and there are some corrections to improve the ZPL value [7, 62]. The most interesting systems in our study have a triplet ground state and involve transitions with the valence or conduction bands, so getting the correct Slater determinant of the excited state is more involved. In the Supplementary Material, we bound this error in the ZPL energy to eV or lower. In all cases, we restricted our study to neutral defects [32].
The formation energy of a defect was calculated by using
where the is the formation energy of a defect , is the defect’s energy, is the energy of a pristine hBN bi- or tri-layer, is the number of a specific atom which is added () or removed () from the defectless hBN system, and is the chemical potential of a atom.
The chemical potentials in Eq.(1) depend on the experimental growth conditions of hBN. Bounds on the values of the chemical potentials of N and B can be established by ab inito thermodynamics method [63]. In this approach, it is assumed that the growth is a process in thermodynamic equilibrium and that the hBN serves as a reservoir of N and B. Therefore, the chemical potentials of B and N are not independent but linked by the formation enthalpy of hBN with respect to bulk B and gaseous ( eV/formula unit) [64, 65]. Hence, two bounds are in order. The N-poor condition in which , and the N-rich condition in which [5]. The chemical potential of carbon is taken as the C-rich bound such that .
To calculate the phonon sideband of photoluminescence spectra, we used the method developed by Alkauskas et al. [66]. In summary, the phonons (assumed to be the same in the ground and excited states) are expanded into generalized coordinates , resonating with the rearrangement upon emission.
where , , and denote atoms, Cartesian coordinates, and the vibrational mode, respectively. and are the positions in excited and relaxed states, respectively. is the atom’s mass. Finally, is the unitary vector for the -th vibrational mode. Thus, we can obtain the partial Huang-Rhys factor , representing the number of phonons of the phonon :
where is the frequency of the phonon -th vibrational mode and is the Planck Constant. From here, the optical spectral function and the phonon sideband can be directly determined [66]. The only free parameter is a thermal-like broadening of , which we set to 25 meV (room temperature).
III Results
All of the defects studied are substitutional C-dimers, with both C atoms in different layers. The cases of CNCB, CNCN and CBCB were studied. In all dimers, the C atom in the bottom layer was kept fixed while we changed the position of the C atom in the top layer, see Fig. 1A. For clarity, we labeled the in-plane position () of one C atom with respect to the other with the letters ‘a’-‘g’. In the configuration ‘a’, both C atoms are on top of each other. The in-plane projection of the C-C distance is in Table 1. The C atoms can belong to adjacent layers () or can be separated by an inner layer (); see Fig. 1B-C. Thus, a given dimer defect can be identified by CXCY–, with B, N.
![Refer to caption](x1.png)
Label | Distance (Å) | Distance C-C (Å) | |
a | |||
b | |||
c | |||
e | |||
f | - | ||
g | - |
Almost all configurations with C defects of the same species, i.e. CNCN and CBCB, have a triplet ground state () regardless of their relative position or whether they belong to adjacent layers or not. As shown in Fig. 2, the triplet is approximately 0.5 eV more stable than the singlet state; these defects will remain spin-active well above room temperature. The only exception is CNCN–b. In this particular defect, both atoms attract each other, moving out of the plane, and they form a covalent bond resembling a sp3 hybridization.
On the other hand, for defects of the type CNCB, the ground state is always a singlet. In Section IV, we discuss why the singlet and triplet spin states are favored for different defect types.
![Refer to caption](x2.png)
Configuration | Hf (eV) | |
N-poor | N-rich | |
CBCB–b | 8.65 | 3.45 |
CBCB–e | 8.66 | 3.46 |
CBCB–a | 8.60 | 3.40 |
CBCB–c | 8.73 | 3.53 |
CBCB–f | 8.59 | 3.39 |
CBCB–g | 8.75 | 3.55 |
CNCN–b | 2.80 | 8.00 |
CNCN–e | 3.67 | 8.87 |
CNCN–a | 3.65 | 8.85 |
CNCN–c | 3.63 | 8.83 |
CNCN–f | 3.66 | 8.86 |
CNCN–g | 3.65 | 8.85 |
CNCB–a | 4.50 | 4.50 |
CNCB–c | 4.47 | 4.47 |
CNCB–b | 4.90 | 4.90 |
CNCB–e | 4.94 | 4.94 |
The formation energy of the CXCY defects is provided in Table 2. Regardless of whether the defects reside in the same or different layers [31], the formation energy is primarily determined by the environment (i.e. the chemical potential). In the most favorable environment (N-rich for CBCB, N-poor for CNCN), eV. Conversely, increases significantly to 8.5-9.0 eV in the least favorable environment. Also, details such as the C–C distance have a minimal effect on the formation energy. The only exception is CNCN–b where a covalent C–C bond develops. Note that this defect has the smallest formation energy among all CNCN defects, which is coherent with a partial covalent bond between both C atoms.
Instead, for CBCN, the formation energies are over eV. This is substantially larger than the dimer C2 (i.e CN, CB adjacent, in the same layer), with eV [32]. As a comparison, topological CNCB defects (i.e. Stone-Wales-like) have a formation energy even higher than the CNCB defects studied here [67], but in these defects decreases quickly with strain. Indeed, biaxial strain can substantially lower the relative formation energy of CB or CN depending on whether it is compressive or tensile [68]. While a large strain is not expected to be present in the bulk, C defects can act as stress valves near grain boundaries [67]. The thermal stability of C defects is determined not only by but also by the energy barriers preventing them from migrating. Babar et al. [67] calculated the barrier energy separating a Stone-Wales C2 dimer from the standard C2 dimer; it surpasses 6 eV, as strong covalent bonds need to be broken. In the case of the CXCY–, the same applies, and we do not expect any migration induced by temperature.
![Refer to caption](x3.png)
The ZPL energy for defects with triplet ground state, CNCN and CBCB, is shown in Fig. 3. When an intermediate layer of hBN separates C atoms, the value of ZPL seems to quickly converge to eV for CNCN and eV for CBCB. When the C atoms are on top of each other or very close, the value of the ZPL decreases. As CNCN–b has a C–C bond, breaking would require a photon energy much larger than the hBN band gap; hence we didn’t calculate its optical excitation. The donor-acceptor-like defects, CNCB– have a ZPL eV, which almost coincides with the ZPL of non-adjacent CNCB when residing in the same layer [25].
![Refer to caption](x4.png)
Two types of photoluminescence spectra were found, see Fig. 4A. The one associated with the defect CBCB–c represents most of the defects studied here; there are high-energy phonon replicas as typically found in C-based defects in hBN. However, there is one exception, CBCB–a, with prominent phonon replicas at 62 meV. In the next section, we will discuss the origin of both types of phonon sidebands found.
IV Discussion
IV.1 Triplet or singlet spin ground state
Almost regardless of distances or whether defects belong to adjacent layers or are separated by an additional hBN layer, defects of the same type (i.e. CBCB, CNCN) have a triplet ground state, well separated in energy from the singlet state ( eV). The same types of defects can have a triplet ground state when embedded in the same layer, but in that case, the singlet and triplet are nearly degenerate [31]. Instead, when the defects belong to a different type (CBCN), the ground state is a singlet.
The relation between the spin state and the type of defect can be understood by looking at the energy of the defect levels within the fundamental gap of hBN. CB has donor-like defect states (close to the conduction band). In contrast, CN is acceptor-like (defect states closer to the valence band). This is illustrated in panels A and D of Fig. 5, where the configurations CBCB–a and CBCN–b are used as examples.
Panel A of Fig. 5 shows the occupied and empty defect levels of a CNCB defect; they are located around CN and CB, respectively (Fig. 5B, C). This directly affects the total spin of the configurations with different types of defects, as it is energetically more favorable for the two extra electrons of the carbon to form a singlet and localize around the CN defect.
On the other hand, defects of the same species - such as CBCB - are degenerate, and when close enough, their wavefunctions are linear combinations of both defect states, as shown in Fig. 5E-F. Also, there is a non-negligible overlap contribution on the B and N atoms along the line joining both C atoms. As the CB states are degenerate, there is an important overlap of the wavefunctions of both occupied levels. This leads to an exchange energy, lowering the energy of the configuration when the total spin is a triplet. In contrast, when the total spin is a singlet, the exchange energy is null, making the energy of the singlet configuration higher than that of the triplet configuration.
![Refer to caption](x5.png)
IV.2 Phonon sideband
When examining the electron-phonon coupling of the optical transition, the relevant phonon replicas are high-energy modes corresponding to bond stretching, typically associated with substitutional C defects [40]. Fig. 4A shows the defect CBCB–c, whose phonon sideband is representative of most of the defects studied here. The optical excitation induces the transfer of one electron from a C atom to the other if the defect is CNCB, or from/to the C atom to the valence/conduction band if the defect is CXCX. Either way, the C–B and/or C–N bonds become weaker or stronger, resulting in a bond stretching: a high-energy localized longitudinal phonon.
However, there is one defect with a markedly different photoluminescence spectrum, CBCB–a, that is, the CB defects are opposite to each other and separated by a N atom from an intermediate layer. This defect features a low-energy phonon replica shifted just meV, see Fig. 4. In this defect, the C atoms are slightly displaced out of the plane (emulating a C–C repulsion) in the ground state, and the phonon associated with such distortion is responsible for the 62 meV phonon replica. In Fig. 4B, C, the phonon with the most relevant partial Huang-Rhys factor is shown for two different dimers (CBCB–a and CBCB–c). This behavior will be discussed in greater detail in the next section. To the best of our knowledge, there are a couple of reports of defects in hBN with a similar low-energy phonon replica, the so-called yellow emitters [69], but these emitters lack spin activity. Although we can discard CBCB–a as yellow emitters, it is likely that the yellow emitters trigger an out-of-plane distortion, explaining their low-energy phonon replicas. The other report of SPEs with a similar PL was obtained in a hBN flake after a H2 plasma treatment [70].
IV.3 Weakly coupled spin pairs
Several reports of SPEs exhibit optically detected magnetic resonance (ODMR) with a very small zero-field splitting (ZFS) in the range of the MHz [20, 71, 72, 73]. Such a small ZFS can be associated with weakly coupled defects [41] or defects with a triplet ground state but with a very small ZFS. Indeed, most CXCX defects studied here are expected to have a small ZFS since both C atoms are separated by nm and the ZFS decays as (in the vacuum). Note that a small ZFS is not incompatible with a large exchange energy, separating the singlet and triplet states. There is a general model aiming to explain the already mentioned ODMR results [42]. It mostly fits with the CNCB–() defects studied here: donor-acceptor pairs, singlet ground state, C–C distance of nm. However, here, we show that weakly interacting triplet states could also be found in hBN multilayers. Particularly, the ODMR measured by Scholten et al [41] effectively behaves as spin 1/2 defects, as the authors demonstrated by Rabi oscillation measurements. Such behavior could occur in defects like the ones from our study, an actual triplet or singlet, but once the exchange energy becomes weaker, that is, when defects are further away. Going closer to the dissociation limit is not only computationally expensive but, more importantly, DFT is known for failing to describe such a limit.
V Conclusions
Substitutional carbon defects are the most widely accepted explanation for the bright light emitters found in hBN. To the best of our knowledge, C-based defects have only been studied when in the same layer.
In this article, we show, by means of DFT calculations, that neutral C defects lying in different layers are also viable candidates to be associated with single emitters. If they form donor-acceptor-like pairs, they behave similarly to when residing in the same hBN layer: they are spinless, have a ZPL energy of eV, and their photoluminescence spectra feature well-defined high-energy phonon replicas.
However, if the two C atoms in different hBN layers are substituting N or B, they have a triplet ground state. It is separated by more than eV from the singlet state, making these defects spin-active above room temperature. Their ZPL can be larger than when the defects form donor-acceptor pairs, eV. Again, these defects present the typical high-energy phonon replicas in their photoluminescence spectra.
When two CB atoms are on top of each other but separated by a hBN interlayer, we calculated a markedly different phonon sideband in the photoluminescence spectrum, with well-defined phonon replicas at meV. They correspond to out-of-plane rearrangement upon excitations. We are unaware of any measurement of such a spectrum for spin-active defects, but the so-called (spinless) yellow emitters have similar phonon replicas and could also correspond to out-of-plane distortions.
VI Competing Interests
The Authors declare no Competing Financial or Non-Financial Interests
VII Data Availability
Input files for the most relevant calculations are provided as supplementary information. The remaining data is available from the corresponding author upon reasonable request.
VIII Author Contributions
I.C. and F.M. ran all the calculations. F.M. conceived the idea. C.C., A.E., and I.C. explained the stability of the triplet state. All authors discussed the results and contributed to the manuscript’s writing.
Acknowledgements.
This work was partially supported by Fondecyt Grants No. 1191353, 1220715, 1220366, 3240387, and 1221512, by the Center for the Development of Nanoscience and Nanotechnology CEDENNA AFB180001, and from Conicyt PIA/Anillo ACT192023. This research was partially supported by the supercomputing infrastructure of the NLHPC (ECM-02).References
- Tran et al. [2016] T. T. Tran, K. Bray, M. J. Ford, M. Toth, and I. Aharonovich, Quantum emission from hexagonal boron nitride monolayers, Nature Nanotechnology 11, 37 (2016).
- Aharonovich et al. [2022] I. Aharonovich, J.-P. Tetienne, and M. Toth, Quantum emitters in hexagonal boron nitride, Nano Letters 22, 9227 (2022).
- Qiu et al. [2024] Z. Qiu, K. Vaklinova, P. Huang, M. Grzeszczyk, K. Watanabe, T. Taniguchi, K. S. Novoselov, J. Lu, and M. Koperski, Atomic and electronic structure of defects in hbn: Enhancing single-defect functionalities, ACS Nano 18, 24035 (2024).
- Mendelson et al. [2021] N. Mendelson, D. Chugh, J. R. Reimers, T. S. Cheng, A. Gottscholl, H. Long, C. J. Mellor, A. Zettl, V. Dyakonov, P. H. Beton, et al., Identifying carbon as the source of visible single-photon emission from hexagonal boron nitride, Nature materials 20, 321 (2021).
- Weston et al. [2018] L. Weston, D. Wickramaratne, M. Mackoit, A. Alkauskas, and C. G. Van de Walle, Native point defects and impurities in hexagonal boron nitride, Phys. Rev. B 97, 214104 (2018).
- Linderälv et al. [2021] C. Linderälv, W. Wieczorek, and P. Erhart, Vibrational signatures for the identification of single-photon emitters in hexagonal boron nitride, Phys. Rev. B 103, 115421 (2021).
- Mackoit-Sinkevičienė et al. [2019] M. Mackoit-Sinkevičienė, M. Maciaszek, C. G. Van de Walle, and A. Alkauskas, Carbon dimer defect as a source of the 4.1 eV luminescence in hexagonal boron nitride, Applied Physics Letters 115, 212101 (2019), https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/1.5124153/13286483/212101_1_online.pdf .
- Du et al. [2015] X. Z. Du, J. Li, J. Y. Lin, and H. X. Jiang, The origin of deep-level impurity transitions in hexagonal boron nitride, Applied Physics Letters 106, 021110 (2015), https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/1.4905908/14088224/021110_1_online.pdf .
- Bourrellier et al. [2016] R. Bourrellier, S. Meuret, A. Tararan, O. Stéphan, M. Kociak, L. H. G. Tizei, and A. Zobelli, Bright uv single photon emission at point defects in h-bn, Nano Letters 16, 4317 (2016).
- Wong et al. [2015] D. Wong, J. J. Velasco, L. Ju, J. Lee, S. Kahn, H.-Z. Tsai, C. Germany, T. Taniguchi, K. Watanabe, A. Zettl, F. Wang, and M. F. Crommie, Nature Nanotechnology 10, 949 (2015).
- Golami et al. [2022] O. Golami, K. Sharman, R. Ghobadi, S. C. Wein, H. Zadeh-Haghighi, C. Gomes da Rocha, D. R. Salahub, and C. Simon, and group theoretical study of properties of a carbon trimer defect in hexagonal boron nitride, Phys. Rev. B 105, 184101 (2022).
- Fischer et al. [2023] M. Fischer, A. Sajid, J. Iles-Smith, A. Hötger, D. I. Miakota, M. K. Svendsen, C. Kastl, S. Canulescu, S. Xiao, M. Wubs, K. S. Thygesen, A. W. Holleitner, and N. Stenger, Combining experiments on luminescent centres in hexagonal boron nitride with the polaron model and ab initio methods towards the identification of their microscopic origin, Nanoscale 15, 14215 (2023).
- Maciaszek and Razinkovas [2024] M. Maciaszek and L. Razinkovas, Blue quantum emitter in hexagonal boron nitride and a carbon chain tetramer: a first-principles study, ACS Applied Nano Materials 7, 18979 (2024).
- Neumann et al. [2023] M. Neumann, X. Wei, L. Morales-Inostroza, S. Song, S.-G. Lee, K. Watanabe, T. Taniguchi, S. Götzinger, and Y. H. Lee, Organic molecules as origin of visible-range single photon emission from hexagonal boron nitride and mica, ACS Nano 17, 11679 (2023).
- Cholsuk et al. [2023] C. Cholsuk, S. Suwanna, and T. Vogl, Comprehensive scheme for identifying defects in solid-state quantum systems, The Journal of Physical Chemistry Letters 14, 6564 (2023).
- Kirchhoff et al. [2022] A. Kirchhoff, T. Deilmann, P. Krüger, and M. Rohlfing, Electronic and optical properties of a hexagonal boron nitride monolayer in its pristine form and with point defects from first principles, Phys. Rev. B 106, 045118 (2022).
- Cassabois et al. [2016] G. Cassabois, P. Valvin, and B. Gil, Hexagonal boron nitride is an indirect bandgap semiconductor, Nature Photonics 10, 262 (2016).
- Román et al. [2021] R. J. P. Román, F. J. R. C. Costa, A. Zobelli, C. Elias, P. Valvin, G. Cassabois, B. Gil, A. Summerfield, T. S. Cheng, C. J. Mellor, P. H. Beton, S. V. Novikov, and L. F. Zagonel, Band gap measurements of monolayer h-BN and insights into carbon-related point defects, 2D Materials 8, 044001 (2021).
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Room temperature coherent control of spin defects in hexagonal boron nitride, Science Advances 7, 10.1126/sciadv.abf3630 (2021).
- Stern et al. [2022] H. L. Stern, Q. Gu, J. Jarman, S. Eizagirre Barker, N. Mendelson, D. Chugh, S. Schott, H. H. Tan, H. Sirringhaus, I. Aharonovich, et al., Room-temperature optically detected magnetic resonance of single defects in hexagonal boron nitride, Nature communications 13, 1 (2022).
- Kianinia et al. [2020] M. Kianinia, S. White, J. E. Fröch, C. Bradac, and I. Aharonovich, Generation of spin defects in hexagonal boron nitride, ACS Photonics 7, 2147 (2020).
- Li et al. [2022] S. Li, A. Pershin, G. Thiering, P. Udvarhelyi, and A. Gali, Ultraviolet quantum emitters in hexagonal boron nitride from carbon clusters, The Journal of Physical Chemistry Letters 13, 3150 (2022).
- Gao et al. [2021] X. Gao, S. Pandey, M. Kianinia, J. Ahn, P. Ju, I. Aharonovich, N. Shivaram, and T. Li, Femtosecond laser writing of spin defects in hexagonal boron nitride, ACS Photonics 8, 994 (2021).
- Acín et al. [2018] A. Acín, I. Bloch, H. Buhrman, T. Calarco, C. Eichler, J. Eisert, D. Esteve, N. Gisin, S. J. Glaser, F. Jelezko, S. Kuhr, M. Lewenstein, M. F. Riedel, P. O. Schmidt, R. Thew, A. Wallraff, I. Walmsley, and F. K. Wilhelm, The quantum technologies roadmap: A european community view, New Journal of Physics 20, 080201 (2018).
- Auburger and Gali [2021] P. Auburger and A. Gali, Towards ab initio identification of paramagnetic substitutional carbon defects in hexagonal boron nitride acting as quantum bits, Phys. Rev. B 104, 075410 (2021).
- Moon et al. [2023] S. Moon, J. Kim, J. Park, S. Im, J. Kim, I. Hwang, and J. K. Kim, Hexagonal boron nitride for next-generation photonics and electronics, Advanced Materials 35, 2204161 (2023).
- Kianinia et al. [2022] M. Kianinia, Z.-Q. Xu, M. Toth, and I. Aharonovich, Quantum emitters in 2d materials: Emitter engineering, photophysics, and integration in photonic nanostructures, Applied Physics Reviews 9, 011306 (2022).
- Çakan et al. [2024] A. Çakan, C. Cholsuk, A. Gale, M. Kianinia, S. Paçal, S. Ateş, I. Aharonovich, M. Toth, and T. Vogl, Quantum optics applications of hexagonal boron nitride defects (2024), arXiv:2410.07712 [quant-ph] .
- Cholsuk et al. [2024] C. Cholsuk, A. Çakan, S. Suwanna, and T. Vogl, Identifying electronic transitions of defects in hexagonal boron nitride for quantum memories, Advanced Optical Materials 12, 2302760 (2024), https://onlinelibrary.wiley.com/doi/pdf/10.1002/adom.202302760 .
- Esmann et al. [2024] M. Esmann, S. C. Wein, and C. Antón-Solanas, Solid-state single-photon sources: Recent advances for novel quantum materials, Advanced Functional Materials 34, 2315936 (2024).
- Pinilla et al. [2023] F. Pinilla, N. Vásquez, I. Chacón, J. R. Maze, C. Cárdenas, and F. Muñoz, Spin-active single photon emitters in hexagonal boron nitride from carbon-based defects, Physica Scripta 98, 095505 (2023).
- Maciaszek et al. [2022] M. Maciaszek, L. Razinkovas, and A. Alkauskas, Thermodynamics of carbon point defects in hexagonal boron nitride, Phys. Rev. Materials 6, 014005 (2022).
- Benedek et al. [2023] Z. Benedek, R. Babar, Á. Ganyecz, T. zilvási, Ö. Legeza, G. Barcza, and V. Ivády, Symmetric carbon tetramers forming spin qubits in hexagonal boron nitride, npj Computational Materials 9, 187 (2023).
- Guo et al. [2021] N.-J. Guo, Y.-Z. Yang, X.-D. Zeng, S. Yu, Y. Meng, Z.-P. Li, Z.-A. Wang, L.-K. Xie, J.-S. Xu, J.-F. Wang, et al., Coherent control of an ultrabright single spin in hexagonal boron nitride at room temperature, arXiv preprint arXiv:2112.06191 (2021).
- Liu et al. [2022] W. Liu, N.-J. Guo, S. Yu, Y. Meng, Z.-P. Li, Y.-Z. Yang, Z.-A. Wang, X.-D. Zeng, L.-K. Xie, Q. Li, J.-F. Wang, J.-S. Xu, Y.-T. Wang, J.-S. Tang, C.-F. Li, and G.-C. Guo, Spin-active defects in hexagonal boron nitride, Materials for Quantum Technology 2, 032002 (2022).
- Qian et al. [2022] C. Qian, V. Villafañe, M. Schalk, G. V. Astakhov, U. Kentsch, M. Helm, P. Soubelet, N. P. Wilson, R. Rizzato, S. Mohr, A. W. Holleitner, D. B. Bucher, A. V. Stier, and J. J. Finley, Unveiling the zero-phonon line of the boron vacancy center by cavity-enhanced emission, Nano Letters 22, 5137 (2022).
- Gottscholl et al. [2020] A. Gottscholl, M. Kianinia, V. Soltamov, S. Orlinskii, G. Mamin, C. Bradac, C. Kasper, K. Krambrock, A. Sperlich, M. Toth, I. Aharonovich, and V. Dyakonov, Initialization and read-out of intrinsic spin defects in a van der waals crystal at room temperature, Nature Materials 19, 540 (2020).
- Huang et al. [2022] P. Huang, M. Grzeszczyk, K. Vaklinova, K. Watanabe, T. Taniguchi, K. S. Novoselov, and M. Koperski, Carbon and vacancy centers in hexagonal boron nitride, Phys. Rev. B 106, 014107 (2022).
- Almohammad et al. [2024] M. Almohammad, Z. Alemoush, J. Li, J. Y. Lin, and H. X. Jiang, Carbon-related donor–acceptor pair transition in the infrared in h-bn, Applied Physics Letters 124, 102106 (2024).
- Jara et al. [2021] C. Jara, T. Rauch, S. Botti, M. A. L. Marques, A. Norambuena, R. Coto, J. E. Castellanos-Águila, J. R. Maze, and F. Munoz, First-principles identification of single photon emitters based on carbon clusters in hexagonal boron nitride, J. P. Chem. A 125, 1325 (2021).
- Scholten et al. [2024] S. C. Scholten, P. Singh, A. J. Healey, I. O. Robertson, G. Haim, C. Tan, D. A. Broadway, L. Wang, H. Abe, T. Ohshima, M. Kianinia, P. Reineck, I. Aharonovich, and J.-P. Tetienne, Multi-species optically addressable spin defects in a van der waals material, Nature Communications 15, 6727 (2024).
- Robertson et al. [2024] I. O. Robertson, B. Whitefield, S. C. Scholten, P. Singh, A. J. Healey, P. Reineck, M. Kianinia, D. A. Broadway, I. Aharonovich, and J.-P. Tetienne, A universal mechanism for optically addressable solid-state spin pairs (2024), arXiv:2407.13148 [cond-mat.mes-hall] .
- Bhang et al. [2021] J. Bhang, H. Ma, D. Yim, G. Galli, and H. Seo, First-principles predictions of out-of-plane group iv and v dimers as high-symmetry, high-spin defects in hexagonal boron nitride, ACS Applied Materials & Interfaces 13, 45768 (2021).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Ab initio molecular dynamics for liquid metals, Phys. Rev. B 47, 558(R) (1993).
- Kresse and Hafner [1994] G. Kresse and J. Hafner, Ab initio molecular-dynamics simulation of the liquid-metal–amorphous-semiconductor transition in germanium, Phys. Rev. B 49, 14251 (1994).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Amblard et al. [2022] D. Amblard, G. D’Avino, I. Duchemin, and X. Blase, Universal polarization energies for defects in monolayer, surface, and bulk hexagonal boron nitride: A finite-size fragments approach, Phys. Rev. Mater. 6, 064008 (2022).
- Winter et al. [2021] M. Winter, M. H. E. Bousquet, D. Jacquemin, I. Duchemin, and X. Blase, Photoluminescent properties of the carbon-dimer defect in hexagonal boron-nitride: A many-body finite-size cluster approach, Phys. Rev. Materials 5, 095201 (2021).
- Badrtdinov et al. [2023] D. I. Badrtdinov, C. Rodriguez-Fernandez, M. Grzeszczyk, Z. Qiu, K. Vaklinova, P. Huang, A. Hampel, K. Watanabe, T. Taniguchi, L. Jiong, M. Potemski, C. E. Dreyer, M. Koperski, and M. Rösner, Dielectric environment sensitivity of carbon centers in hexagonal boron nitride, Small 19, 2300144 (2023).
- Pinilla et al. [2024] F. Pinilla, W. A. Muriel, J. Cabezas-Escares, I. Chacón, C. Cárdenas, and F. Muñoz, Manipulating the stacking in two-dimensional hexagonal boron nitride bilayers: Implications for defect-based single photon emitters, ACS Applied Nano Materials 7, 6039 (2024).
- Gilbert et al. [2019] S. M. Gilbert, T. Pham, M. Dogan, S. Oh, B. Shevitski, G. Schumm, S. Liu, P. Ercius, S. Aloni, M. L. Cohen, and A. Zettl, Alternative stacking sequences in hexagonal boron nitride, 2D Materials 6, 021006 (2019).
- Herath et al. [2019] U. Herath, P. Tavadze, X. He, E. Bousquet, S. Singh, F. Muñoz, and A. H. Romero, Pyprocar: A python library for electronic structure pre/post-processing, Comput. Phys. Commun. 251, 107080 (2019).
- Lang et al. [2024] L. Lang, P. Tavadze, A. Tellez, E. Bousquet, H. Xu, F. Muñoz, N. Vasquez, U. Herath, and A. H. Romero, Expanding pyprocar for new features, maintainability, and reliability, Computer Physics Communications 297, 109063 (2024).
- Momma and Izumi [2011] K. Momma and F. Izumi, Vesta 3 for three-dimensional visualization of crystal, volumetric and morphology data, Journal of applied crystallography 44, 1272 (2011).
- Sun et al. [2015] J. Sun, A. Ruzsinszky, and J. P. Perdew, Strongly constrained and appropriately normed semilocal density functional, Phys. Rev. Lett. 115, 036402 (2015).
- Rauch et al. [2021] T. Rauch, F. Munoz, M. A. L. Marques, and S. Botti, Defect levels from scan and mbj meta-gga exchange-correlation potentials, Phys. Rev. B 104, 064105 (2021).
- Krukau et al. [2006] A. V. Krukau, O. A. Vydrov, A. F. Izmaylov, and G. E. Scuseria, Influence of the exchange screening parameter on the performance of screened hybrid functionals, J. Chem. Phys. 125, 224106 (2006).
- Deák et al. [2010] P. Deák, B. Aradi, T. Frauenheim, E. Janzén, and A. Gali, Accurate defect levels obtained from the hse06 range-separated hybrid functional, Phys. Rev. B 81, 153203 (2010).
- Yang and Ayers [2024] W. Yang and P. W. Ayers, Foundation for the scf approach in density functional theory (2024), arXiv:2403.04604 [physics.chem-ph] .
- Jin et al. [2021] Y. Jin, M. Govoni, G. Wolfowicz, S. E. Sullivan, F. J. Heremans, D. D. Awschalom, and G. Galli, Photoluminescence spectra of point defects in semiconductors: Validation of first-principles calculations, Phys. Rev. Mater. 5, 084603 (2021).
- Iwanski et al. [2024] J. Iwanski, K. P. Korona, M. Tokarczyk, G. Kowalski, A. K. Dabrowska, P. Tatarczak, I. Rogala, M. Bilska, M. Wojcik, S. Kret, A. Reszka, B. J. Kowalski, S. Li, A. Pershin, A. Gali, J. Binder, and A. Wysmolek, Revealing polytypism in 2d boron nitride with uv photoluminescence (2024), arXiv:2405.19126 [cond-mat.mtrl-sci] .
- Reuter and Scheffler [2001] K. Reuter and M. Scheffler, Composition, structure, and stability ofruo2(110)as a function of oxygen pressure, Physical Review B 65, 10.1103/PhysRevB.65.035406 (2001).
- Tomaszkiewicz [2002] I. Tomaszkiewicz, The enthalpy of formation of hexagonal boron nitride, Polish Journal of Chemistry Vol. 76 / nr 6, 891–899 (2002).
- Jr. [1998] M. C. Jr., NIST-JANAF thermochemical tables, monograph 9, Vol. 9 (American Chemical Society, 1998) pp. 1–1951.
- Alkauskas et al. [2014] A. Alkauskas, B. B. Buckley, D. D. Awschalom, and C. G. Van de Walle, First-principles theory of the luminescence lineshape for the triplet transition in diamond nv centres, New J. Phys. 16, 073026 (2014).
- Babar et al. [2024] R. Babar, I. A. Abrikosov, G. Barcza, and V. Ivády, Carbon-contaminated topological defects in hbn: a potential new class of single photon emitters, arXiv preprint arXiv:2403.00755 (2024).
- Santra et al. [2024] P. Santra, S. Ghaderzadeh, M. Ghorbani-Asl, H.-P. Komsa, E. Besley, and A. V. Krasheninnikov, Strain-modulated defect engineering of two-dimensional materials, npj 2D Materials and Applications 8, 33 (2024).
- Kumar et al. [2023] A. Kumar, C. Cholsuk, A. Zand, M. N. Mishuk, T. Matthes, F. Eilenberger, S. Suwanna, and T. Vogl, Localized creation of yellow single photon emitting carbon complexes in hexagonal boron nitride, APL Materials 11, 071108 (2023).
- Chen et al. [2021] Y. Chen, C. Li, S. White, M. Nonahal, Z.-Q. Xu, K. Watanabe, T. Taniguchi, M. Toth, T. T. Tran, and I. Aharonovich, Generation of high-density quantum emitters in high-quality, exfoliated hexagonal boron nitride, ACS Applied Materials & Interfaces 13, 47283 (2021).
- Yang et al. [2023] Y.-Z. Yang, T.-X. Zhu, Z.-P. Li, X.-D. Zeng, N.-J. Guo, S. Yu, Y. Meng, Z.-A. Wang, L.-K. Xie, Z.-Q. Zhou, Q. Li, J.-S. Xu, X.-Y. Gao, W. Liu, Y.-T. Wang, J.-S. Tang, C.-F. Li, and G.-C. Guo, Laser direct writing of visible spin defects in hexagonal boron nitride for applications in spin-based technologies, ACS Applied Nano Materials 6, 6407 (2023).
- Guo et al. [2023] N.-J. Guo, S. Li, W. Liu, Y.-Z. Yang, X.-D. Zeng, S. Yu, Y. Meng, Z.-P. Li, Z.-A. Wang, L.-K. Xie, R.-C. Ge, J.-F. Wang, Q. Li, J.-S. Xu, Y.-T. Wang, J.-S. Tang, A. Gali, C.-F. Li, and G.-C. Guo, Coherent control of an ultrabright single spin in hexagonal boron nitride at room temperature, Nature Communications 14, 2893 (2023).
- Chejanovsky et al. [2021] N. Chejanovsky, A. Mukherjee, J. Geng, Y.-C. Chen, Y. Kim, A. Denisenko, A. Finkler, T. Taniguchi, K. Watanabe, D. B. R. Dasari, et al., Single-spin resonance in a van der waals embedded paramagnetic defect, Nature materials 20, 1079 (2021).