### v2

# Optical Properties and Spin States of Inter-layer Carbon Defect Pairs in Hexagonal Boron Nitride: A First-Principles Study

Abstract
Substitutional carbon defects in hexagonal boron nitride (hBN) are prominent single photon emitters (SPEs), and their potential for spin activity () is particularly intriguing. While studies have largely focused on intra-layer defects, we employ density functional theory (DFT) to investigate inter-layer dimers of identical carbon species (CXCX). We demonstrate that these CXCX pairs can exhibit a stable triplet spin state at room temperature when closely spaced (e.g., within 3.5-7.1 Å) across hBN layers. As their separation increases beyond this range (e.g., Å), they transition into weakly interacting pairs, characterized by singlet-triplet degeneracy. This regime is predicted to result in a very small zero-field splitting for the triplet manifold, offering a potential explanation for certain optically detected magnetic resonance (ODMR) signals. The zero-phonon line (ZPL) energy of these inter-layer CXCX pairs is found to be practically monochromatic and within the visible range. Furthermore, we identify specific CBCB inter-layer configurations exhibiting atypical low-energy phonon replicas due to out-of-plane vibrational coupling, a finding that may clarify the vibronic structure of other hBN emitters, such as the ‘yellow emitters’.
I Introduction
Hexagonal boron nitride (hBN) is a layered material that can be exfoliated down to the monolayer. It has garnered attention for hosting single photon emitters [1, 2, 3]. Many of these defects have been associated with substitutional carbon defects [1, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], even though other options have been proposed [14, 15, 16]. Due to its large bandgap, close to 6 eV, defects incorporated within the crystal can introduce new energy levels within the gap [17, 18]. Consequently, electronic transitions give rise to SPEs exhibiting characteristics such as bright photons, a commendable coherence time [19], even at room temperature, and in some cases, they are spin-active (triplet or higher spin state) defects [20, 21, 22, 23]. These properties are crucial for quantum technologies in communication, computation, metrology, sensing, imaging, etc. [24, 25, 26, 27, 28, 29, 30]. Therefore, the search for these types of defects is relevant, considering the scarcity of substitutional spin-active carbon defects, such as CNCN [31], (CN)3CB and (CB)3CN [32, 33]. It is worth mentioning that vacancy-based SPEs in hBN are an active research field, as they often are spin-active [34, 35, 36, 37].
Substitutional C defects behave as donors (CB), acceptors (CN), donor-acceptor pairs, or more complex arrangements [38, 39]. Each C defect contributes a single electron and one defect state (spin up and down) within the hBN band gap. The 4.1 eV emission line has been identified as two adjacent CNCB defects [7], often denoted as C2. If the same CN–CB pair is in the same layer but separated by some distance, the emission energy decreases to the visible range [40, 25]. The ground state is likely spinless if the number of CN and CB is the same. Otherwise, the system can have a non-zero net spin. There are experimental findings of emissions attributed to C defects, even though the actual C defect has not been identified [20, 41]. One possible explanation, fitting with the available results, is weakly coupled defects [42].
This article aims to investigate SPEs when forming a C dimer but residing in different hBN layers. Such a possibility has not been addressed before, even though defects comprising an interstitial C atom have been considered [43]. The paper is structured as follows: Section II details the methods employed for the calculations, Section III presents our main results, including the SPE’s spin state, likely to be a triplet, and their photoluminescence spectra. Finally, Section IV explains the large energy difference between triplet and singlet states for some dimers, the phonon sidebands shape, and we relate our findings to experimental results on spin-active SPEs.
II Methodology
II.1 Computational Methods
Density functional theory (DFT) calculations were performed using the Vienna ab initio simulation package (VASP) [44, 45, 46, 47]. We utilized a hexagonal supercell of hBN with -point sampling. Defects in adjacent layers were modeled using an hBN bilayer, while an hBN trilayer was employed for defects in non-adjacent layers. These multi-layer models aim to approximate bulk hBN; additional layers are expected to influence the zero-phonon line (ZPL) energy through (i) enhanced dielectric screening [48, 49, 50] and (ii) modified electrostatic potentials [51], with these influences diminishing as layer thickness increases towards the bulk limit. The AA’ stacking (B and N atoms on top of N and B, respectively), characteristic of most hBN samples [52], was adopted.
The kinetic energy cutoff was set to 500 eV. PyProcar [53, 54] was used for the analysis of the results, and VESTA for visualization [55].
The SCAN exchange-correlation (XC) functional [56] was employed for structural relaxation, phonon calculations, and optical transition energies. The suitability of SCAN for defect levels is supported when the host’s band gap is reasonably described [57], as is the case here (SCAN-predicted hBN gap: 5.0 eV; experimental value eV). While the Heyd-Scuseria-Ernzerhof (HSE06) functional[58] is a common alternative, its results are highly sensitive to dielectric screening, which is significantly reduced and anisotropic in 2D materials like hBN. This sensitivity is pronounced for donor-acceptor defects with strong electrostatic interactions. Consequently, applying HSE06 to our bilayer and trilayer models, which inherently possess different screening characteristics, yields results whose direct comparison or extrapolation to bulk behavior is complex. Nevertheless, HSE06 optical transition energies, calculated using the SCAN-relaxed geometries (a common practice where less demanding functionals provide reliable structures [59]), are reported in the Supplementary Material. On the other hand, the functional of Perdew-Burke-Ernzerhof [60] failed to describe the ground state of the defects at large distances, providing a spurious triplet state. Even worse, when using the PBE’s wavefunctions as seeds for more expensive calculations, the error is propagated.
To study the excited state of each configuration, the SCF method was used [61, 62]. This method has a problem of correctly describing the excited state (i.e. mixed spin states are employed instead of well-defined spins), and there are some corrections to improve the ZPL value [7, 63]. The most interesting systems in our study have a triplet ground state and involve transitions with the valence or conduction bands, so getting the correct Slater determinant of the excited state is too involved to be practical. In the Supplementary Material, we bound this error in the ZPL energy to eV or lower. In all cases, we restricted our study to neutral defects [32]. Calculation of the single-photon emission rates would involve the computation of electric dipole transition moments between the involved excited and ground electronic states, which is beyond the scope of the current work.
The formation energy of a defect was calculated by using
where the is the formation energy of a defect , is the defect’s energy, is the energy of a pristine hBN bi- or tri-layer, is the number of a specific atom which is added () or removed () from the defectless hBN system, and is the chemical potential of a atom.
The chemical potentials in Eq. (1) depend on the experimental growth conditions of hBN. Bounds on the values of the chemical potentials of N and B can be established by ab inito thermodynamics method [64]. In this approach, it is assumed that the growth is a process in thermodynamic equilibrium and that the hBN serves as a reservoir of N and B. Therefore, the chemical potentials of B and N are not independent but linked by the formation enthalpy of hBN with respect to bulk B and gaseous ( eV/formula unit) [65, 66]. Hence, two bounds are in order. The N-poor condition in which , and the N-rich condition in which [5]. The chemical potential of carbon is taken as the C-rich bound such that .
To calculate the phonon sideband of photoluminescence spectra, we used the method developed by Alkauskas et al. [67]. In summary, the phonons (assumed to be the same in the ground and excited states) are expanded into generalized coordinates , resonating with the rearrangement upon emission.
where , , and denote atoms, Cartesian coordinates, and the vibrational mode, respectively. and are the positions in excited and relaxed states, respectively. is the atom’s mass. Finally, is the unitary vector for the -th vibrational mode. Thus, we can obtain the partial Huang-Rhys factor , representing the number of phonons of the phonon :
where is the frequency of the phonon -th vibrational mode and is the Planck constant. The total Huang-Rhys factor is . The photoluminescence spectral lineshape was then constructed using the generating function formalism, which builds the vibronic progression based on , the set of , and [67]. Finally, the calculated stick spectrum (ZPL and phonon replicas) was convoluted with a Gaussian function with a Full Width at Half Maximum (FWHM) of 25 meV. This phenomenological broadening accounts for homogeneous lifetime effects, inhomogeneous broadening, and approximates thermal effects at room temperature, allowing for comparison with experimental spectra.
From here, the optical spectral function and the phonon sideband can be directly determined [67]. The only free parameter is a thermal-like broadening of , which we set to 25 meV (room temperature).
III Results
All of the defects studied are substitutional C-dimers, with both C atoms in different layers. The cases of CNCB, CNCN and CBCB were studied. In all dimers, the C atom in the bottom layer was kept fixed while we changed the position of the C atom in the top layer, see Fig. 1A. For clarity, we labeled the in-plane position () of one C atom with respect to the other by the in-plane projection of their distance, taking the nearest neighbor distance as ‘1’ (i.e. is equivalent to 1.45 Å), as found in a similar context for labeling C-defects [25]. To convert to times the lattice parameter, it is needed to multiply it by . The conversion of to Å is given in Table 1. The C atoms can belong to adjacent layers () or can be separated by an inner layer (); see Fig. 1B-C. Thus, a given dimer defect can be identified by CXCY–, with B, N.
| Distance (nearest neighbors) | Distance (Å) | Distance C-C (Å) | |
Defects of the type CNCB are consistently found in our calculations to exhibit a singlet ground state (see Fig. 2). This observation aligns with established expectations: CN sites, characterized by defect energy levels close to the valence band maximum, act as acceptors. Conversely, CB sites are understood to behave as donors. Consequently, a charge transfer from CB to CN is anticipated, irrespective of their separation, leading to effectively charged C and C centers. In the dissociation limit, as direct quantum mechanical interactions (such as exchange coupling) between the two carbon atoms vanish, their distinct electronic natures –CB as a donor and CN as an acceptor– remain the dominant factor. Overall, the system will naturally remain a singlet, consistent with our findings.
In contrast, carbon defects of the same species, i.e., CNCN and CBCB, generally exhibit a stable triplet ground state when the defects are relatively close. This tendency is typically observed for separations up to (an in-plane distance of Å) and (a vertical separation of two layers, resulting in a total distance of approximately 7Å), as detailed in Fig. 2. However, for CBCB defects in adjacent layers (), the spin state at shorter distances can be more sensitive to the specific geometric configuration.
The spin state of the defect pair is intrinsically linked to the symmetry of the two-electron spatial wavefunction: a singlet state is associated with a spatially symmetric wavefunction, while a triplet state corresponds to a spatially anti-symmetric wavefunction. This spatial symmetry dictates the charge distribution of the pair; for instance, an anti-symmetric wavefunction (triplet) possesses a nodal plane or region of zero electron density at the midpoint between the defects, whereas a symmetric wavefunction (singlet) typically has finite electron density there. This difference in charge distribution influences the electrostatic interaction with the polar hBN lattice, a mechanism previously invoked to explain the pronounced stability of the triplet state for certain CNCN configurations within the same hBN layer (h=0) [31].
As the defects are moved further apart, the significance of the wavefunction’s symmetry at the midpoint diminishes because the overall wavefunction amplitude in this central region approaches zero. In this dissociation limit, the C atoms evolve into weakly interacting S=1/2 paramagnetic centers [42]. Consequently, the exchange interaction between them vanishes, leading to the singlet and triplet states becoming effectively degenerate, as anticipated for two distant, uncorrelated S=1/2 spins.
A notable exception to the general triplet preference for nearby CNCN defects is the CNCN–(1,1) configuration. In this case, the two CN atoms experience a strong mutual attraction, causing them to displace significantly from the hBN plane and form a direct covalent C-C bond. This bonding scenario, which locally resembles sp3 hybridization, results in a singlet ground state, i.e. consistent with the spin-pairing typically observed upon formation of a strong covalent bond.
| Configuration | Hf (eV) | |
| N-poor | N-rich | |
| CBCB– | 9.05 | 3.85 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.07 | 3.87 |
| CBCB– | 9.07 | 3.87 |
| CBCB– | 9.06 | 3.86 |
| CBCB– | 9.06 | 3.86 |
| CBCB– | 9.05 | 3.85 |
| CBCB– | 9.06 | 3.86 |
| CNCN– | 2.88 | 8.08 |
| CNCN– | 3.47 | 8.67 |
| CNCN– | 3.44 | 8.64 |
| CNCN– | 3.45 | 8.65 |
| CNCN– | 3.45 | 8.65 |
| CNCN– | 3.42 | 8.62 |
| CNCN– | 3.40 | 8.60 |
| CNCN– | 3.40 | 8.60 |
| CNCN– | 3.41 | 8.61 |
| CNCB– | 4.71 | 4.71 |
| CNCB– | 4.75 | 4.75 |
| CNCB– | 4.84 | 4.84 |
| CNCB– | 5.13 | 5.13 |
| CNCB– | 5.14 | 5.14 |
| CNCB– | 5.15 | 5.15 |
| CNCB– | 5.17 | 5.17 |
The formation energies () of the investigated CXCY defects are presented in Table 2. For CBCB and CNCN defects, is primarily determined by the chemical environment (i.e., the chemical potential of nitrogen). Under their most favorable conditions –N-rich for CBCB and N-poor for CNCN– the formation energies for CBCB configurations are found to be in the narrow range of 3.85-3.88 eV, while for CNCN configurations they range from 2.88 eV to 3.47 eV. Conversely, in their least favorable environments –N-poor for CBCB and N-rich for CNCN– increases substantially, to approximately 9.05-9.08 eV for CBCB and 8.08-8.67 eV for CNCN.
Generally, for these homonuclear pairs (CBCB and CNCN), variations in the specific C–C distance () or interlayer arrangement () have a relatively minor impact on (typically less than 0.05 eV), especially when compared to the large shifts induced by the chemical environment. A prominent exception to this trend is the CNCN–(1,1) configuration. It exhibits the lowest formation energy among all CNCN defects under both N-poor (2.88 eV) and N-rich (8.08 eV) conditions. This enhanced stability, representing a reduction of approximately 0.5-0.6 eV compared to other CNCN configurations at , is consistent with the formation of a stabilizing covalent C–C bond, as discussed in relation to its structure and spin state.
For the heteronuclear CNCB defects, the calculated formation energies (Table 2) range from approximately 4.71 eV to 5.17 eV. These formation energies are identical under N-poor and N-rich conditions because the creation of a CNCB pair involves substituting one B and one N atom with two C atoms; thus, the terms related to and in the formation energy expression effectively cancel each other. The formation energies for the inter-layer CNCB configurations investigated here (all with or , and not forming direct covalent C-C bonds between the pair) are substantially larger than the eV reported for an adjacent CN-CB pair within the same hBN layer () [32]. The significantly lower energy of this intra-layer C2 dimer is likely attributable to the formation of a direct C–C covalent bond, a strong stabilizing effect similar to that observed for the CNCN–(1,1) configuration in our study.
For broader context, other types of carbon defects, such as topological CNCB defects (e.g., Stone-Wales-like structures), can exhibit even higher formation energies than the substitutional CNCB pairs in Table 2, although their can be significantly reduced by mechanical strain [68]. Indeed, strain engineering through biaxial stress has been shown to modulate the formation energies of individual CB or CN defects [69]. While large, uniform strains are not typically expected in bulk hBN, localized stresses, for instance near grain boundaries, could influence defect formation and stability, with carbon defects potentially acting to relieve such stresses [68].
The overall thermal stability and persistence of these C defects in hBN are governed not only by their formation energies but also critically by the kinetic barriers against their migration or transformation. For instance, Babar et al. [68] calculated a high energy barrier (over 6 eV) for the transformation between a Stone-Wales C2 dimer and a standard substitutional C2 dimer, highlighting that C-C bond rearrangements can be significantly hindered. While specific migration and transformation pathways for the CXCY–() configurations in our study have not been computed, similar considerations of strong covalent bond stability suggest that their diffusion and spontaneous restructuring are also likely to be impeded by substantial energy barriers, limiting thermally induced migration under typical experimental conditions.
To further underscore the expected kinetic stability of these C–C pairs, we can estimate the energy cost associated with a specific decomposition mechanism. Consider a CBCB pair, for instance, CBCB–(). One possible route for its partial decomposition involves one of the carbon atoms transitioning from its boron substitutional site (leaving it as VB) into an interstitial position (Ci), while the other carbon atom remains as a CB. The products of such a process would be an isolated CB defect, a carbon interstitial Ci, and a boron vacancy VB. We can estimate the formation energy of this dissociated state, , using the formation energies of these individual neutral defects.
These values render the proposed decomposition pathway highly improbable, even at high temperature (i.e., a barrier over 10 eV). While is about one half of the formation energy of a CB dimer, the interstitial and the vacancy have a much higher , which is expected, as in both defects, strong covalent bonds become dangling bonds. Beyond such decomposition mechanisms, local structural transformations could also occur. For example, an in-plane rotation of a CB-CN bond pair, a process analogous to a step in Stone-Wales defect formation, would constitute another pathway for atomic rearrangement. Again, covalent bonds need to be broken and reformed, and a very high barrier is expected for such processes in hBN [68].
A similar rationale would apply to the decomposition of CNCN pairs, provided the formation energies of the analogous dissociated products (e.g., CN, Ci, VN) result in a similarly high overall energy cost. This analysis, while specific to one decomposition channel, supports the general expectation that transforming these substitutional carbon pairs into defect configurations involving interstitials and vacancies is kinetically strongly hindered. Overall, the robustness of the covalent network in hBN implies that most pathways for significant structural alteration or migration of these carbon defect pairs will involve high activation energies.
The ZPL energy for the defects studied is shown in Fig. 3, as calculated with the SCAN XC functional. In the supplementary material, Fig. S1 shows the values calculated with HSE06 (see the Computational Methods). The donor-acceptor-like dimers have a ZPL energy decreasing with the C–C distance. As the individual defects are charged, their electrostatic interaction is considerable even at long distances. The emission should be in the green-orange range of the visible spectrum. The use of HSE06 is inconsistent for these heteronuclear defects, as we used a different thickness for , effectively changing the dielectric constant.
Defect dimers of the same type, CNCN and CBCB, have a ZPL practically independent of the C–C distance (the smallest distance considered is over 3.5 Å). For CBCB defects, the ZPL is close to eV, and for CNCN it is close to eV, see Fig. 3. To have a reference value, with the HSE06 hybrid functional (Fig. 1 in the supplementary information), the ZPL energies of CBCB and CNCN are 2.1 and 2.4 eV, respectively. Again, the values are almost independent of the C–C distance. It is worth mentioning that, as there is no strong electrostatic interaction, this time the values of the HSE06 functional are not heavily affected by changing the dielectric constant. It is well documented that DFT has an inherent problem in determining accurately the band gap of materials [70]. This is particularly true for defect states [71]. Then, additional information, such as the vibrational spectra, is useful to determine the defect nature against experimental data [40].
The CNCN–(1,1) configuration, which is characterized by the formation of a direct C-C covalent bond, was considered a distinct case. Its optical electronic transition involving this strong C-C bond (e.g., ) should occur at energies larger than the band gap of hBN. For this reason, its ZPL associated with such deep defect states was not computed.
Two types of photoluminescence spectra were found, see Fig. 4A. The one associated with the defect CBCB– represents most of the defects studied here. Figures S2–S4 in the supplementary material show the photoluminescence spectra of all the studied defects emitting within the visible range. These high-energy phonon replicas are typically found in C-based defects in hBN [40, 1]. They correspond to (in-plane) bond stretching involving the C atoms and their closest neighbors, see Fig. 4B. While so far, only defects and defect clusters within a single layer had been considered, the mechanism behind the activation of this phonon mode should apply regardless of the exact location of both C atoms: one electron jumps between defect atoms, or from/to a defect level and the host bands, strengthening or weakening some bonds and causing the in-plane stretching.
However, there are a couple of exceptions, CBCB– (shown in Fig. 4C), and CBCB– with prominent phonon replicas of meV. In these cases, an out-of-plane phonon mode resonates with the transition. In the next section, we will discuss the origin of both types of phonon sidebands found.
IV Discussion
IV.1 Triplet or singlet spin ground state
The relation between the spin state and the type of defect pair can be understood by examining the energy and nature of the defect levels within the fundamental band gap of hBN. As established, CB sites introduce donor-like defect states, positioned closer to the conduction band minimum. In contrast, CN sites create acceptor-like defect states, situated nearer to the valence band maximum. This distinction is illustrated schematically in panels A (for a CNCB type) and D (for a CBCB type) of Fig. 5, using the CBCN– and CBCB– configurations as respective examples.
Detailed electronic band structures for many of the defect configurations studied are provided in Figures S5-S7 of the Supplementary Material. For cases where singlet and triplet states of a defect pair were found to be nearly degenerate in energy, only the band structure for the triplet state is presented, as the singlet counterpart yields almost identical results. Generally, the donor/acceptor picture for CB/CN holds as long as the carbon defects maintain their sp2 hybridization within the hBN plane. A notable exception is the CNCN– configuration (see Fig. S7 in the Supplementary Material), which undergoes significant structural rearrangement to form a C-C bond with sp3-like character. For this defect, a single, non-degenerate defect-related band appears within the hBN band gap, with other associated levels located deep within the valence band.
For heteronuclear CNCB defect pairs, the inherent donor (CB) and acceptor (CN) nature of the constituent defects leads to charge transfer from CB to CN, resulting in C and C ionic states. This is depicted in Fig. 5A, where the occupied defect levels (associated with C) are shown lower in the gap, and empty levels (associated with C) are higher. The corresponding wavefunctions (Fig. 5B,C) illustrate the spatial localization of these occupied and empty states primarily on the CN and CB sites, respectively. This charge transfer results in a doubly occupied acceptor state localized on the CN site (forming C) and an emptied donor site (forming C). As these resulting ionic centers (C and C) are effectively non-magnetic due to their spin-paired electronic configurations, the CNCB defect pair consequently exhibits a singlet ground state.
In contrast, for homonuclear defect pairs, such as CBCB (or CNCN), both components are of the same type (e.g., both donors for CBCB). For these homonuclear pairs, the identical nature of the two defects means their respective unperturbed defect states are intrinsically degenerate. When these defects interact at sufficiently short distances, this degeneracy is lifted, and new combined states are formed which can be described as linear combinations of the original defect states, as illustrated for CBCB in Fig. 5D-F. For these systems, each C atom contributes an electron to the defect-related manifold. The significant overlap between the wavefunctions of these occupied levels gives rise to an exchange interaction. This interaction typically stabilizes the triplet state (parallel spins), where the spatial wavefunction is antisymmetric, relative to the singlet state (antiparallel spins), which has a symmetric spatial wavefunction, thus making the triplet state the ground state at shorter distances. As the separation between the identical defects increases, the wavefunction overlap and the consequent exchange interaction diminish. In this limit, the energy difference between the singlet and triplet states vanishes, leading to their degeneracy as the system effectively becomes two weakly interacting centers.
IV.2 Phonon sideband
When examining the electron-phonon coupling of the optical transition, the relevant phonon replicas are high-energy modes corresponding to bond stretching, typically associated with substitutional C defects [40]. Fig. 4A shows the defect CBCB–, whose phonon sideband is representative of most of the defects studied here, Figures S2-S4 in the supplementary material have the photoluminescence spectra for all remaining configurations. The optical excitation induces the transfer of one electron from a C atom to the other if the defect is CNCB, or from/to the C atom to the valence/conduction band if the defect is CXCX. Either way, the C–B and/or C–N bonds become weaker or stronger, resulting in a bond stretching: a high-energy localized longitudinal phonon.
However, there are two defects with a markedly different photoluminescence spectra, CBCB– and CBCB–. In the former, the CB defects are opposite to each other and separated by a N atom from an intermediate layer. This defect features a low-energy phonon replica shifted just meV, see Fig. 4A. In this defect, the C atoms are slightly displaced out of the plane (emulating a C–C repulsion) in the ground state, and the phonon associated with such distortion is responsible for the 62 meV phonon replica. A similar scenario occurs in CBCB–. Fig. 4B, C, shows the phonon with the most relevant partial Huang-Rhys factor , for two different dimers, CBCB– and CBCB–, while in the former the electronic excitation induces are rearrangement restricted to the defect plane, in the later the dominant phonon modes are out of plane.
To the best of our knowledge, there are a couple of reports of defects in hBN with a similar low-energy phonon replica, the so-called yellow emitters [72], but these emitters lack spin activity. Although we can discard CBCB– and CBCB– as yellow emitters, it is likely that the yellow emitters trigger an out-of-plane distortion, explaining their low-energy phonon replicas. The other report of SPEs with a similar PL was obtained in a hBN flake after a H2 plasma treatment [73].
IV.3 Weakly coupled spin pairs
Several reports of SPEs in hBN exhibit optically detected magnetic resonance (ODMR) with very small zero-field splittings (ZFS) in the MHz range [20, 74, 75, 76]. Such small ZFS values can arise either from weakly coupled, physically distinct defects [41] or from individual defects possessing a triplet ground state with an intrinsically small ZFS. Our calculated CXCX defects are candidates for both scenarios: when the two carbon atoms are separated by 0.4-0.7 nm, the ZFS within their triplet state is expected to be small, primarily due to the decay of the magnetic dipole-dipole interaction. It is important to note that a small ZFS within the triplet manifold is not incompatible with a large exchange energy () that separates the singlet and triplet states, a situation relevant for our CXCX defects at shorter C-C distances (see Fig. 2). At larger separations, as the defects become spin degenerate.
At this point, it is informative to estimate the ZFS of the CXCX defects where the constituent carbon atoms are separated by distances in the range of 0.4-1.0 nm; the magnetic dipole-dipole interaction is expected to be a primary contributor. Using the point-dipole approximation (, with MHz nm3, see Section V from the supplementary material), we estimate the ZFS parameter to range from approximately 78 MHz (for nm) up to 1.22 GHz (for nm). For a specific configuration like CBCB– with an inter-carbon distance of 0.67 nm, this estimation yields a value of approximately 260 MHz. These values, particularly at the larger end of the separation range, are substantially smaller than the multi-GHz ZFS typically observed for compact spin-1 defects like V in hBN [77], and fall within a range that could be consistent with ODMR experiments reporting MHz-scale features.
A general model by Robertson et al. [42] proposes that some of the experimentally observed ODMR signals arise from pairs of weakly coupled defects. While our CNCB– defects share some characteristics with the systems considered by this model, such as C-C distances around nm, their established singlet ground state (due to charge transfer) makes them unlikely candidates to be these pairs themselves.
This distinction motivates exploring other defect types. Here, we propose that our CXCX homonuclear defect pairs, provide a compelling alternative origin for such ODMR signatures. As discussed, when these CXCX defects are sufficiently far apart, the exchange energy () diminishes, and the system behaves as two weakly interacting centers. The resulting state would exhibit a very small ZFS, as argued above. Such a system could indeed produce ODMR signals that appear to arise from ‘effectively defects,’ consistent with the experimental findings of Scholten et al. [41], who observed this characteristic behavior through Rabi oscillation measurements. Thus, the CXCX pairs at in hBN multilayers naturally reconcile the picture of weakly interacting spins with an underlying triplet state possessing a minimal ZFS.
IV.4 Suitability or usefulness as SPEs
Overall, the studied defects present a mix of typical behaviors for in-plane defects and new features. Heteronuclear CBCN defect pairs are predicted to behave as the typical bright SPEs usually found in hBN [1]. They have a typical phonon sideband with high-energy replicas (160-190 meV), and their emission covers the entire visible range, also they usually do not produce a measurable ODMR signal [28]. All these characteristics suggest a large family of related defects, as proposed by different arrangements of substitutions C defects acting as donor-acceptor systems [40, 25]. Our contribution shows that in such a family, C defects residing in different layers are also candidates for SPEs.
Instead, the homonuclear defects CBCB, CNCN show some interesting features for SPEs, (i) often they are spin-active at room temperature (i.e., a triplet) and (ii) regardless of details, each series (CBCB or CNCN) has a very similar ZPL, converging to an asymptotic limit. Both features are not present in C dimers within the same monolayer [31] and are interesting for SPEs. On one hand, most of the already known spin-active defects in hBN are based on vacancies [35], which share many favorable and detrimental aspects with well-studied SPEs such as the NV-center in diamond (e.g., triplet ground state, ZFS in the GHz range, broad emission line). Instead, most defects spin-active CBCB, CNCN defects would have some interesting properties as SPEs: a bright emission, a narrow emission line, and well-separated phonon replicas. But, their ZFS would be non-constant and range from tens to hundreds of MHz, which could make the integration with current technology difficult.
There are two particular defect configurations, CBCB– and CBCB–, with a phonon sideband different, to the best of our knowledge, from other substitutional C-based SPEs, characterized by prominent meV replicas from out-of-plane modes. If this distinct vibronic structure leads to an effectively broader overall emission profile where the ZPL is less dominant or more difficult to spectrally isolate, it could hinder applications requiring high photon indistinguishability and long coherence times, as these rely heavily on the properties of the ZPL photons.
Arguably, the most significant problem of SPEs in hBN is the difficulty of making monochromatic SPEs within the visible. So far, there are two irradiation protocols to achieve monodisperse emitters: the so-called blue [78] and the yellow [72] emitters. In this respect, very specific defect geometries -such as CBCB–- are unlikely to be deterministically created. Instead, there are irradiation protocols aimed at producing defects [79]. Such protocols could be adapted [80] to produce pairs of within less than a nm. In the presence of carbon contamination [4], these defects could turn into CBCB dimers as those studied here, some of them with a triplet ground state.
Are the pairs of C defects considered here SPEs at all? For strongly coupled CXCX pairs forming a triplet state, or for CNCB singlet pairs, the pair acts as a single quantum emitter. In the limit of weakly interacting CXCX pairs, while the two centers are nearly independent, single-photon emission would still be expected if the optical cycle involves defined joint states of the pair. Definitive confirmation of single-photon emission characteristics relies on measurements. As stated before, measurements of hBN defects of yet undetermined atomic structure reveal a very small ZFS tens of MHz or even smaller. These reports routinely show the dip associated with single-photon emission. This is a strong hint that well-separated defects can behave as SPEs. Particularly, one article [20] shows a defect in hBN with a ZFS of 50 MHz, together with a photoluminescence spectrum featuring the typical high-energy phonon replicas (180 meV) of C-based defects in hBN and a ZPL of eV. All these measurement points to CNCN defects of the type studied here, perhaps with a larger C–C distance (1.1-1.2 nm).
Even though we didn’t calculate the emission rates of the defects studied here, it is relevant to ask if these weakly interacting pairs have an emission rate relevant for applications (or even detection). We can not give a definitive answer. However, the experiments mentioned above suggest a large enough emission rate, as the single defects provide a clear signal. For the donor-acceptor-like pairs, the charge transfer should imply a large dipole transition moment and a large emission rate. Experimentally, their “brightness is the highest reported for a quantum emitter in the visible spectral range” [1].
V Conclusions
Substitutional carbon defects are the most widely accepted explanation for the bright light emitters found in hBN. To the best of our knowledge, C-based defects have only been studied when in the same layer.
In this article, we show, by means of DFT calculations, that neutral C defects lying in different layers are also viable candidates to be associated with single emitters. If they form donor-acceptor-like pairs, they behave similarly to when residing in the same hBN layer: they are spinless, have a ZPL energy of eV, and their photoluminescence spectra feature well-defined high-energy phonon replicas.
However, if the two C atoms in different hBN layers are substituting the same species, N or B, they likely have a triplet ground state for closer distances Å. It is separated by more than eV from the singlet state, making these defects spin-active above room temperature. Their ZPL is almost independent of the C-C distance (i.e. nearly monochromatic), according to SCAN-based calculations, their emission is within the red. Calculations with the HSE06 functional shift the emission energy to the yellow-to-green range. Regardless of the actual value of the ZPL, the nearly monochromatic emission holds. Most of these kinds of defect dimers present the typical high-energy phonon replicas in their photoluminescence spectra.
As the distance of the mononuclear defects (CBCB or CNCN) increases, they become practically spin degenerate, or ‘weakly interacting spin pairs’ [42]. In either case, a robust triplet or a spin degenerate state, the defects are well-separated and their zero field splitting should be within the MHz range, which is consistent with some optically detected magnetic resonance (ODMR) experiments [41, 20, 74, 75, 76].
When two CB atoms are on top of each other but separated by a hBN interlayer or lying as close as possible in adjacent layers -CBCB– and CBCB-- their photoluminescence spectrum has a markedly different phonon sideband, with well-defined phonon replicas at meV. They correspond to out-of-plane rearrangement upon excitations. We are unaware of any measurement of such a spectrum for spin-active defects, but the so-called (spinless) yellow emitters [72] have similar phonon replicas and could also correspond to out-of-plane distortions.
VI Competing Interests
The Authors declare no Competing Financial or Non-Financial Interests
VII Data Availability
Input files for the most relevant calculations are provided as supplementary information. The remaining data is available from the corresponding author upon reasonable request.
VIII Author Contributions
I.C. and F.M. ran all the calculations. F.M. conceived the idea. C.C., A.E., and I.C. explained the stability of the triplet state. All authors discussed the results and contributed to the manuscript’s writing.
Acknowledgements.
This work was partially supported by Fondecyt Grants No. 1191353, 1220715, 1220366, 3240387, and 1221512, by the Center for the Development of Nanoscience and Nanotechnology CEDENNA AFB180001, and from Conicyt PIA/Anillo ACT192023. This research was partially supported by the supercomputing infrastructure of the NLHPC (ECM-02).References
- Tran et al. [2016] T. T. Tran, K. Bray, M. J. Ford, M. Toth, and I. Aharonovich, Quantum emission from hexagonal boron nitride monolayers, Nature Nanotechnology 11, 37 (2016).
- Aharonovich et al. [2022] I. Aharonovich, J.-P. Tetienne, and M. Toth, Quantum emitters in hexagonal boron nitride, Nano Letters 22, 9227 (2022).
- Qiu et al. [2024] Z. Qiu, K. Vaklinova, P. Huang, M. Grzeszczyk, K. Watanabe, T. Taniguchi, K. S. Novoselov, J. Lu, and M. Koperski, Atomic and electronic structure of defects in hbn: Enhancing single-defect functionalities, ACS Nano 18, 24035 (2024).
- Mendelson et al. [2021] N. Mendelson, D. Chugh, J. R. Reimers, T. S. Cheng, A. Gottscholl, H. Long, C. J. Mellor, A. Zettl, V. Dyakonov, P. H. Beton, et al., Identifying carbon as the source of visible single-photon emission from hexagonal boron nitride, Nature materials 20, 321 (2021).
- Weston et al. [2018] L. Weston, D. Wickramaratne, M. Mackoit, A. Alkauskas, and C. G. Van de Walle, Native point defects and impurities in hexagonal boron nitride, Phys. Rev. B 97, 214104 (2018).
- Linderälv et al. [2021] C. Linderälv, W. Wieczorek, and P. Erhart, Vibrational signatures for the identification of single-photon emitters in hexagonal boron nitride, Phys. Rev. B 103, 115421 (2021).
- Mackoit-Sinkevičienė et al. [2019] M. Mackoit-Sinkevičienė, M. Maciaszek, C. G. Van de Walle, and A. Alkauskas, Carbon dimer defect as a source of the 4.1 eV luminescence in hexagonal boron nitride, Applied Physics Letters 115, 212101 (2019), https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/1.5124153/13286483/212101_1_online.pdf .
- Du et al. [2015] X. Z. Du, J. Li, J. Y. Lin, and H. X. Jiang, The origin of deep-level impurity transitions in hexagonal boron nitride, Applied Physics Letters 106, 021110 (2015), https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/1.4905908/14088224/021110_1_online.pdf .
- Bourrellier et al. [2016] R. Bourrellier, S. Meuret, A. Tararan, O. Stéphan, M. Kociak, L. H. G. Tizei, and A. Zobelli, Bright uv single photon emission at point defects in h-bn, Nano Letters 16, 4317 (2016).
- Wong et al. [2015] D. Wong, J. J. Velasco, L. Ju, J. Lee, S. Kahn, H.-Z. Tsai, C. Germany, T. Taniguchi, K. Watanabe, A. Zettl, F. Wang, and M. F. Crommie, Nature Nanotechnology 10, 949 (2015).
- Golami et al. [2022] O. Golami, K. Sharman, R. Ghobadi, S. C. Wein, H. Zadeh-Haghighi, C. Gomes da Rocha, D. R. Salahub, and C. Simon, and group theoretical study of properties of a carbon trimer defect in hexagonal boron nitride, Phys. Rev. B 105, 184101 (2022).
- Fischer et al. [2023] M. Fischer, A. Sajid, J. Iles-Smith, A. Hötger, D. I. Miakota, M. K. Svendsen, C. Kastl, S. Canulescu, S. Xiao, M. Wubs, K. S. Thygesen, A. W. Holleitner, and N. Stenger, Combining experiments on luminescent centres in hexagonal boron nitride with the polaron model and ab initio methods towards the identification of their microscopic origin, Nanoscale 15, 14215 (2023).
- Maciaszek and Razinkovas [2024] M. Maciaszek and L. Razinkovas, Blue quantum emitter in hexagonal boron nitride and a carbon chain tetramer: a first-principles study, ACS Applied Nano Materials 7, 18979 (2024).
- Neumann et al. [2023] M. Neumann, X. Wei, L. Morales-Inostroza, S. Song, S.-G. Lee, K. Watanabe, T. Taniguchi, S. Götzinger, and Y. H. Lee, Organic molecules as origin of visible-range single photon emission from hexagonal boron nitride and mica, ACS Nano 17, 11679 (2023).
- Cholsuk et al. [2023] C. Cholsuk, S. Suwanna, and T. Vogl, Comprehensive scheme for identifying defects in solid-state quantum systems, The Journal of Physical Chemistry Letters 14, 6564 (2023).
- Kirchhoff et al. [2022] A. Kirchhoff, T. Deilmann, P. Krüger, and M. Rohlfing, Electronic and optical properties of a hexagonal boron nitride monolayer in its pristine form and with point defects from first principles, Phys. Rev. B 106, 045118 (2022).
- Cassabois et al. [2016] G. Cassabois, P. Valvin, and B. Gil, Hexagonal boron nitride is an indirect bandgap semiconductor, Nature Photonics 10, 262 (2016).
- Román et al. [2021] R. J. P. Román, F. J. R. C. Costa, A. Zobelli, C. Elias, P. Valvin, G. Cassabois, B. Gil, A. Summerfield, T. S. Cheng, C. J. Mellor, P. H. Beton, S. V. Novikov, and L. F. Zagonel, Band gap measurements of monolayer h-BN and insights into carbon-related point defects, 2D Materials 8, 044001 (2021).
- Gottscholl et al. [2021a] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Room temperature coherent control of spin defects in hexagonal boron nitride, Science Advances 7, 10.1126/sciadv.abf3630 (2021a).
- Stern et al. [2022] H. L. Stern, Q. Gu, J. Jarman, S. Eizagirre Barker, N. Mendelson, D. Chugh, S. Schott, H. H. Tan, H. Sirringhaus, I. Aharonovich, et al., Room-temperature optically detected magnetic resonance of single defects in hexagonal boron nitride, Nature communications 13, 1 (2022).
- Kianinia et al. [2020] M. Kianinia, S. White, J. E. Fröch, C. Bradac, and I. Aharonovich, Generation of spin defects in hexagonal boron nitride, ACS Photonics 7, 2147 (2020).
- Li et al. [2022] S. Li, A. Pershin, G. Thiering, P. Udvarhelyi, and A. Gali, Ultraviolet quantum emitters in hexagonal boron nitride from carbon clusters, The Journal of Physical Chemistry Letters 13, 3150 (2022).
- Gao et al. [2021] X. Gao, S. Pandey, M. Kianinia, J. Ahn, P. Ju, I. Aharonovich, N. Shivaram, and T. Li, Femtosecond laser writing of spin defects in hexagonal boron nitride, ACS Photonics 8, 994 (2021).
- Acín et al. [2018] A. Acín, I. Bloch, H. Buhrman, T. Calarco, C. Eichler, J. Eisert, D. Esteve, N. Gisin, S. J. Glaser, F. Jelezko, S. Kuhr, M. Lewenstein, M. F. Riedel, P. O. Schmidt, R. Thew, A. Wallraff, I. Walmsley, and F. K. Wilhelm, The quantum technologies roadmap: A european community view, New Journal of Physics 20, 080201 (2018).
- Auburger and Gali [2021] P. Auburger and A. Gali, Towards ab initio identification of paramagnetic substitutional carbon defects in hexagonal boron nitride acting as quantum bits, Phys. Rev. B 104, 075410 (2021).
- Moon et al. [2023] S. Moon, J. Kim, J. Park, S. Im, J. Kim, I. Hwang, and J. K. Kim, Hexagonal boron nitride for next-generation photonics and electronics, Advanced Materials 35, 2204161 (2023).
- Kianinia et al. [2022] M. Kianinia, Z.-Q. Xu, M. Toth, and I. Aharonovich, Quantum emitters in 2d materials: Emitter engineering, photophysics, and integration in photonic nanostructures, Applied Physics Reviews 9, 011306 (2022).
- Çakan et al. [2024] A. Çakan, C. Cholsuk, A. Gale, M. Kianinia, S. Paçal, S. Ateş, I. Aharonovich, M. Toth, and T. Vogl, Quantum optics applications of hexagonal boron nitride defects (2024), arXiv:2410.07712 [quant-ph] .
- Cholsuk et al. [2024] C. Cholsuk, A. Çakan, S. Suwanna, and T. Vogl, Identifying electronic transitions of defects in hexagonal boron nitride for quantum memories, Advanced Optical Materials 12, 2302760 (2024), https://onlinelibrary.wiley.com/doi/pdf/10.1002/adom.202302760 .
- Esmann et al. [2024] M. Esmann, S. C. Wein, and C. Antón-Solanas, Solid-state single-photon sources: Recent advances for novel quantum materials, Advanced Functional Materials 34, 2315936 (2024).
- Pinilla et al. [2023] F. Pinilla, N. Vásquez, I. Chacón, J. R. Maze, C. Cárdenas, and F. Muñoz, Spin-active single photon emitters in hexagonal boron nitride from carbon-based defects, Physica Scripta 98, 095505 (2023).
- Maciaszek et al. [2022] M. Maciaszek, L. Razinkovas, and A. Alkauskas, Thermodynamics of carbon point defects in hexagonal boron nitride, Phys. Rev. Materials 6, 014005 (2022).
- Benedek et al. [2023] Z. Benedek, R. Babar, Á. Ganyecz, T. zilvási, Ö. Legeza, G. Barcza, and V. Ivády, Symmetric carbon tetramers forming spin qubits in hexagonal boron nitride, npj Computational Materials 9, 187 (2023).
- Guo et al. [2021] N.-J. Guo, Y.-Z. Yang, X.-D. Zeng, S. Yu, Y. Meng, Z.-P. Li, Z.-A. Wang, L.-K. Xie, J.-S. Xu, J.-F. Wang, et al., Coherent control of an ultrabright single spin in hexagonal boron nitride at room temperature, arXiv preprint arXiv:2112.06191 (2021).
- Liu et al. [2022] W. Liu, N.-J. Guo, S. Yu, Y. Meng, Z.-P. Li, Y.-Z. Yang, Z.-A. Wang, X.-D. Zeng, L.-K. Xie, Q. Li, J.-F. Wang, J.-S. Xu, Y.-T. Wang, J.-S. Tang, C.-F. Li, and G.-C. Guo, Spin-active defects in hexagonal boron nitride, Materials for Quantum Technology 2, 032002 (2022).
- Qian et al. [2022] C. Qian, V. Villafañe, M. Schalk, G. V. Astakhov, U. Kentsch, M. Helm, P. Soubelet, N. P. Wilson, R. Rizzato, S. Mohr, A. W. Holleitner, D. B. Bucher, A. V. Stier, and J. J. Finley, Unveiling the zero-phonon line of the boron vacancy center by cavity-enhanced emission, Nano Letters 22, 5137 (2022).
- Gottscholl et al. [2020] A. Gottscholl, M. Kianinia, V. Soltamov, S. Orlinskii, G. Mamin, C. Bradac, C. Kasper, K. Krambrock, A. Sperlich, M. Toth, I. Aharonovich, and V. Dyakonov, Initialization and read-out of intrinsic spin defects in a van der waals crystal at room temperature, Nature Materials 19, 540 (2020).
- Huang et al. [2022] P. Huang, M. Grzeszczyk, K. Vaklinova, K. Watanabe, T. Taniguchi, K. S. Novoselov, and M. Koperski, Carbon and vacancy centers in hexagonal boron nitride, Phys. Rev. B 106, 014107 (2022).
- Almohammad et al. [2024] M. Almohammad, Z. Alemoush, J. Li, J. Y. Lin, and H. X. Jiang, Carbon-related donor–acceptor pair transition in the infrared in h-bn, Applied Physics Letters 124, 102106 (2024).
- Jara et al. [2021] C. Jara, T. Rauch, S. Botti, M. A. L. Marques, A. Norambuena, R. Coto, J. E. Castellanos-Águila, J. R. Maze, and F. Munoz, First-principles identification of single photon emitters based on carbon clusters in hexagonal boron nitride, J. P. Chem. A 125, 1325 (2021).
- Scholten et al. [2024] S. C. Scholten, P. Singh, A. J. Healey, I. O. Robertson, G. Haim, C. Tan, D. A. Broadway, L. Wang, H. Abe, T. Ohshima, M. Kianinia, P. Reineck, I. Aharonovich, and J.-P. Tetienne, Multi-species optically addressable spin defects in a van der waals material, Nature Communications 15, 6727 (2024).
- Robertson et al. [2024] I. O. Robertson, B. Whitefield, S. C. Scholten, P. Singh, A. J. Healey, P. Reineck, M. Kianinia, D. A. Broadway, I. Aharonovich, and J.-P. Tetienne, A universal mechanism for optically addressable solid-state spin pairs (2024), arXiv:2407.13148 [cond-mat.mes-hall] .
- Bhang et al. [2021] J. Bhang, H. Ma, D. Yim, G. Galli, and H. Seo, First-principles predictions of out-of-plane group iv and v dimers as high-symmetry, high-spin defects in hexagonal boron nitride, ACS Applied Materials & Interfaces 13, 45768 (2021).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Ab initio molecular dynamics for liquid metals, Phys. Rev. B 47, 558(R) (1993).
- Kresse and Hafner [1994] G. Kresse and J. Hafner, Ab initio molecular-dynamics simulation of the liquid-metal–amorphous-semiconductor transition in germanium, Phys. Rev. B 49, 14251 (1994).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Amblard et al. [2022] D. Amblard, G. D’Avino, I. Duchemin, and X. Blase, Universal polarization energies for defects in monolayer, surface, and bulk hexagonal boron nitride: A finite-size fragments approach, Phys. Rev. Mater. 6, 064008 (2022).
- Winter et al. [2021] M. Winter, M. H. E. Bousquet, D. Jacquemin, I. Duchemin, and X. Blase, Photoluminescent properties of the carbon-dimer defect in hexagonal boron-nitride: A many-body finite-size cluster approach, Phys. Rev. Materials 5, 095201 (2021).
- Badrtdinov et al. [2023] D. I. Badrtdinov, C. Rodriguez-Fernandez, M. Grzeszczyk, Z. Qiu, K. Vaklinova, P. Huang, A. Hampel, K. Watanabe, T. Taniguchi, L. Jiong, M. Potemski, C. E. Dreyer, M. Koperski, and M. Rösner, Dielectric environment sensitivity of carbon centers in hexagonal boron nitride, Small 19, 2300144 (2023).
- Pinilla et al. [2024] F. Pinilla, W. A. Muriel, J. Cabezas-Escares, I. Chacón, C. Cárdenas, and F. Muñoz, Manipulating the stacking in two-dimensional hexagonal boron nitride bilayers: Implications for defect-based single photon emitters, ACS Applied Nano Materials 7, 6039 (2024).
- Gilbert et al. [2019] S. M. Gilbert, T. Pham, M. Dogan, S. Oh, B. Shevitski, G. Schumm, S. Liu, P. Ercius, S. Aloni, M. L. Cohen, and A. Zettl, Alternative stacking sequences in hexagonal boron nitride, 2D Materials 6, 021006 (2019).
- Herath et al. [2019] U. Herath, P. Tavadze, X. He, E. Bousquet, S. Singh, F. Muñoz, and A. H. Romero, Pyprocar: A python library for electronic structure pre/post-processing, Comput. Phys. Commun. 251, 107080 (2019).
- Lang et al. [2024] L. Lang, P. Tavadze, A. Tellez, E. Bousquet, H. Xu, F. Muñoz, N. Vasquez, U. Herath, and A. H. Romero, Expanding pyprocar for new features, maintainability, and reliability, Computer Physics Communications 297, 109063 (2024).
- Momma and Izumi [2011] K. Momma and F. Izumi, Vesta 3 for three-dimensional visualization of crystal, volumetric and morphology data, Journal of applied crystallography 44, 1272 (2011).
- Sun et al. [2015] J. Sun, A. Ruzsinszky, and J. P. Perdew, Strongly constrained and appropriately normed semilocal density functional, Phys. Rev. Lett. 115, 036402 (2015).
- Rauch et al. [2021] T. Rauch, F. Munoz, M. A. L. Marques, and S. Botti, Defect levels from scan and mbj meta-gga exchange-correlation potentials, Phys. Rev. B 104, 064105 (2021).
- Krukau et al. [2006] A. V. Krukau, O. A. Vydrov, A. F. Izmaylov, and G. E. Scuseria, Influence of the exchange screening parameter on the performance of screened hybrid functionals, J. Chem. Phys. 125, 224106 (2006).
- Deák et al. [2010] P. Deák, B. Aradi, T. Frauenheim, E. Janzén, and A. Gali, Accurate defect levels obtained from the hse06 range-separated hybrid functional, Phys. Rev. B 81, 153203 (2010).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Yang and Ayers [2024] W. Yang and P. W. Ayers, Foundation for the scf approach in density functional theory (2024), arXiv:2403.04604 [physics.chem-ph] .
- Jin et al. [2021] Y. Jin, M. Govoni, G. Wolfowicz, S. E. Sullivan, F. J. Heremans, D. D. Awschalom, and G. Galli, Photoluminescence spectra of point defects in semiconductors: Validation of first-principles calculations, Phys. Rev. Mater. 5, 084603 (2021).
- Iwanski et al. [2024] J. Iwanski, K. P. Korona, M. Tokarczyk, G. Kowalski, A. K. Dabrowska, P. Tatarczak, I. Rogala, M. Bilska, M. Wojcik, S. Kret, A. Reszka, B. J. Kowalski, S. Li, A. Pershin, A. Gali, J. Binder, and A. Wysmolek, Revealing polytypism in 2d boron nitride with uv photoluminescence (2024), arXiv:2405.19126 [cond-mat.mtrl-sci] .
- Reuter and Scheffler [2001] K. Reuter and M. Scheffler, Composition, structure, and stability ofruo2(110)as a function of oxygen pressure, Physical Review B 65, 10.1103/PhysRevB.65.035406 (2001).
- Tomaszkiewicz [2002] I. Tomaszkiewicz, The enthalpy of formation of hexagonal boron nitride, Polish Journal of Chemistry Vol. 76 / nr 6, 891–899 (2002).
- Jr. [1998] M. C. Jr., NIST-JANAF thermochemical tables, monograph 9, Vol. 9 (American Chemical Society, 1998) pp. 1–1951.
- Alkauskas et al. [2014] A. Alkauskas, B. B. Buckley, D. D. Awschalom, and C. G. Van de Walle, First-principles theory of the luminescence lineshape for the triplet transition in diamond nv centres, New J. Phys. 16, 073026 (2014).
- Babar et al. [2024] R. Babar, I. A. Abrikosov, G. Barcza, and V. Ivády, Carbon-contaminated topological defects in hbn: a potential new class of single photon emitters, arXiv preprint arXiv:2403.00755 (2024).
- Santra et al. [2024] P. Santra, S. Ghaderzadeh, M. Ghorbani-Asl, H.-P. Komsa, E. Besley, and A. V. Krasheninnikov, Strain-modulated defect engineering of two-dimensional materials, npj 2D Materials and Applications 8, 33 (2024).
- Perdew [1985] J. P. Perdew, Density functional theory and the band gap problem, International Journal of Quantum Chemistry 28, 497 (1985).
- Reimers et al. [2018] J. R. Reimers, A. Sajid, R. Kobayashi, and M. J. Ford, Understanding and calibrating density-functional-theory calculations describing the energy and spectroscopy of defect sites in hexagonal boron nitride, Journal of Chemical Theory and Computation 14, 1602 (2018).
- Kumar et al. [2023] A. Kumar, C. Cholsuk, A. Zand, M. N. Mishuk, T. Matthes, F. Eilenberger, S. Suwanna, and T. Vogl, Localized creation of yellow single photon emitting carbon complexes in hexagonal boron nitride, APL Materials 11, 071108 (2023).
- Chen et al. [2021] Y. Chen, C. Li, S. White, M. Nonahal, Z.-Q. Xu, K. Watanabe, T. Taniguchi, M. Toth, T. T. Tran, and I. Aharonovich, Generation of high-density quantum emitters in high-quality, exfoliated hexagonal boron nitride, ACS Applied Materials & Interfaces 13, 47283 (2021).
- Yang et al. [2023] Y.-Z. Yang, T.-X. Zhu, Z.-P. Li, X.-D. Zeng, N.-J. Guo, S. Yu, Y. Meng, Z.-A. Wang, L.-K. Xie, Z.-Q. Zhou, Q. Li, J.-S. Xu, X.-Y. Gao, W. Liu, Y.-T. Wang, J.-S. Tang, C.-F. Li, and G.-C. Guo, Laser direct writing of visible spin defects in hexagonal boron nitride for applications in spin-based technologies, ACS Applied Nano Materials 6, 6407 (2023).
- Guo et al. [2023] N.-J. Guo, S. Li, W. Liu, Y.-Z. Yang, X.-D. Zeng, S. Yu, Y. Meng, Z.-P. Li, Z.-A. Wang, L.-K. Xie, R.-C. Ge, J.-F. Wang, Q. Li, J.-S. Xu, Y.-T. Wang, J.-S. Tang, A. Gali, C.-F. Li, and G.-C. Guo, Coherent control of an ultrabright single spin in hexagonal boron nitride at room temperature, Nature Communications 14, 2893 (2023).
- Chejanovsky et al. [2021] N. Chejanovsky, A. Mukherjee, J. Geng, Y.-C. Chen, Y. Kim, A. Denisenko, A. Finkler, T. Taniguchi, K. Watanabe, D. B. R. Dasari, et al., Single-spin resonance in a van der waals embedded paramagnetic defect, Nature materials 20, 1079 (2021).
- Gottscholl et al. [2021b] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hbn as promising temperature, pressure and magnetic field quantum sensors, Nature Communications 12, 4480 (2021b).
- Gale et al. [0] A. Gale, C. Li, Y. Chen, K. Watanabe, T. Taniguchi, I. Aharonovich, and M. Toth, Site-specific fabrication of blue quantum emitters in hexagonal boron nitride, ACS Photonics 0, null (0).
- Carbone et al. [2025] A. Carbone, I. D. Breev, J. Figueiredo, S. Kretschmer, L. Geilen, A. B. Mhenni, J. Arceri, A. V. Krasheninnikov, M. Wubs, A. W. Holleitner, A. Huck, C. Kastl, and N. Stenger, Quantifying the creation of negatively charged boron vacancies in he-ion irradiated hexagonal boron nitride (2025), arXiv:2501.18481 [cond-mat.mes-hall] .
- Ren et al. [2025] F. Ren, Z. Xu, and Y. Wu, Optimization of carbon irradiation parameters for creating spin defects in hexagonal boron nitride, Nanotechnology and Precision Engineering 8, 033014 (2025), https://pubs.aip.org/tu/npe/article-pdf/doi/10.1063/10.0036119/20472607/033014_1_10.0036119.pdf .